steph. wants to calculate the volume of her beach ball. the diameter of the beach ball is 24 inches. waht is its volume?
a. 6912 pi cu. in.
b. 2304 pi cu. in
c 576 pi cu in
d. 196 pi cu in

Answer by ewatrrr(24785)

 
Hi,
diameter of the beach ball is 24 inches, raius = 12
V = (4/3)pir^3 = (4/3)pi*(12in)^3 = 2304pi in^3